OpenStudy (anonymous):

@matthew1286 i need ur help

OpenStudy (matthew1286):

with what im no that good at math lol but ill see if i can help

OpenStudy (anonymous):

@jessie13

OpenStudy (anonymous):

do u knw math

OpenStudy (jessie13):

okay

OpenStudy (anonymous):

so u knwo math

OpenStudy (jessie13):

yes

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Simon has a garden that is 10 feet wide. The perimeter of the garden is 50 feet. What is the area of Simon’s garden?

OpenStudy (anonymous):

i got 150 is that right

OpenStudy (rhystic):

yeah

OpenStudy (jessie13):

yes that is right! that is what i got!

OpenStudy (anonymous):

i got 3 more can u help me

OpenStudy (jessie13):

sure

OpenStudy (anonymous):

The rug in Mario’s room is 6 feet by 6 feet. He wants to buy a new rug that is double the length of the old rug. How will the area of the rug change? A. The area of the new rug will be the same as the area of the old rug. B. The area of the new rug will be 4 times the area of the old rug. C. The area of the new rug will be half the area of the old rug. D. The area of the new rug will be twice the area of the old rug.

OpenStudy (anonymous):

i got D for the answer

OpenStudy (jessie13):

yes D is right!

OpenStudy (matthew1286):

i told u d dude

OpenStudy (matthew1286):

i told u i was sure it was d

OpenStudy (anonymous):

i was just double checking

OpenStudy (matthew1286):

ok well i said it was d u didnt say anything

OpenStudy (matthew1286):

good bye

OpenStudy (anonymous):

@Smileyside what do u think

OpenStudy (jessie13):

i got 200!

OpenStudy (anonymous):

u sure

OpenStudy (matthew1286):

i was thinking 200 but i was never fully sure

OpenStudy (anonymous):

ok i got another one ready matthew and the rest of u guys

OpenStudy (jessie13):

k

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

help does anyone know this

OpenStudy (rhystic):

find the area of the larger square and then of the small squares subtract the small squares value from the large square value and divide by 4 yielding the amount of tiles needed.

OpenStudy (anonymous):

thats what i did i got 130

OpenStudy (rhystic):

your math is wrong? The area of the big square is 8*8 = 64 ; area of 1 small square is 2*2 = 4 but there are four of them so 4*4 = 16 work it from there and see what you get.

OpenStudy (anonymous):

so i do what doi need to add or subtract or multiply

OpenStudy (rhystic):

you need to subtract the small square area from the larger to get the remaining area to be filled up with white tiles. After you have the remaning area you have to divide by 4 which is the area of a small tile to see how many small tiles will actually be needed to take up the left over space.

OpenStudy (anonymous):

can u just give me the equation to solve

OpenStudy (rhystic):

64-16 = ? -> ?/4 = answer

OpenStudy (anonymous):

i got 54

OpenStudy (anonymous):

is that right